# Polarimetric compressed sensing with hollow, self-assembled diffractive films

### v1

@@ 2 Results
When samples are ordered, the simulated resonances are narrow and located at an angle of incidence $\theta_{i}\approx\pm 15-20^{\circ}$ depending on the rotation of the sample. Experimentally, this TE mode is strongest in ordered PANI-HS samples and disappears as samples are more random, evident from the specular transmission and reflection measurements with a lock-amplifier (See Supporting information Measurements and Simulations, Fig. S7). Across different experimental batch samples, TM-polarized modes are inconsistently observed, and these variations could relate to the grain mesostructure, fiber nucleation statistics, or PANI doping. However, over many batches and over years in ambient conditions, the sample TE lattice mode $\theta_{i}\approx\pm 15-20^{\circ}$ remain strong and are consistent from batch to batch. The periodic modulations of ${\rm S}_{2}$ associated with the TE-polarized lattice modes occur with angular periods of $\sim 40^{\circ}$ and encode the incident state of polarization.
Notably, the random PANI-HS sample offers a higher degree of incoherence compared to the ordered one with broadened $k$-space angular modes, less-distinct fringes, and broadened modulation of the far-field polarization. This enhanced incoherence for the same input information indicates that the random sample achieves a higher degree of compression than the ordered one. Even so, for each $\theta_{i}$ there remain many $\theta_{o}$ far-field locations that are “dark”. Therefore, further broadening via the smoothing or blurring of signals would further increase the degree of compressive sensing.
@@
Figure 7: Multi-parameter estimation of incident beam polarization and pointing. (a) Illustration of 4-encoder-detector system for simulation with 150-degree field-of-view. The detector separation for this case is $\Delta\theta_{o}=37.5^{\circ}$. (b) Colormaps illustrating difference LP-signals given the geometry in (a) for raw experimental data of ordered sample (left) and 5-degree window-blurred sample (right). Compression ratios as a function of detector spacing with (c) raw data and (d) blurred data for ordered (blue) and random (green) samples. The markers show the mean values and shaded areas mark the upper and lower bounds from 15 different test simulations with polynomial fits between points. Inset: the pointing and polarization ($S_{2},S_{3}$) prediction errors. Dotted vertical lines mark Nyquist sampling of the pointing or when $\Delta\theta_{o}=$ Pointing MAE $\approx 20^{\circ}$. Best compression ratios are achieved with Nyquist sampling twice the frequency of the interference fringes or $\Delta\theta_{o}\approx 4^{\circ}$.
@@
From this dataset that includes the full-Stokes characteristics at each angle, we consider the sparse “detector measurements” from two crossed LP detectors aligned ±$\pm 45^{\circ}$ from the sensing-plane. Figure 7(b) shows colormaps of the LP-difference detected signals from each encoder given the geometry in Fig. 7(a) for RCP and LCP. The LP-difference detected signals are the logarithms of the intensity signals multipled by the ${\rm S}_{2}$ measurements. The raw experimental data from ordered samples appear noisy and have narrow, angularly-sharp features; in fact, a significant amount of the information lies in the low-power measurements that are noisy. We smooth measurements with a Savitzky-Golay filter [Fig. 7(b,right)]. This filter is analogous to the signal integration that is expected to be present in nocturnal insects and under low-light conditions 26, 27 or with micro-saccadic head movements.
where $\Delta\theta_{o}$ is the average detector spacing, $\varepsilon_{\rm Pointing}$, $\varepsilon_{\rm S_{2}}$, and $\varepsilon_{\rm S_{3}}$ are the SNN prediction mean absolute error (MAE) associated with pointing, ${\rm S}_{2}$ and ${\rm S}_{3}$, respectively. The factors of $\pi$ in Eq. 9 refer to the area of the equatorial-plane unit circle of the Poincare sphere. Eq. 9 is an approximation, since we do not predict ${\rm S}_{0}$ and our inputs are ${\rm S}_{2}$, ${\rm S}_{3}=\pm 1$ while we consider a regression model where ${\rm S}_{2}$ and ${\rm S}_{3}$ are independent variables. For the purpose of this investigation, Eq. 9 is a reliable measure since ${C_{\rm R}}$ decreases when the polarization prediction error increases. Moreover, for a solid beam spot, the pointing MAE is approximately half of the detector spacing $\varepsilon_{\rm Pointing}\sim\Delta\theta_{o}/2$ and when $\varepsilon_{\rm S_{2}},\varepsilon_{\rm S_{3}}=0$, Eq. 9 yields ${C_{\rm R}}=2$. This factor of two is expected since the full-Stokes polarization characterization typically requires four separate measurements while our encoded approach employs only two.
The SNN predictions indicate two notable transitions, denoted by the compression ratio ${C_{\rm R}}$ [Fig. 7(c-d)] and the range of the error predictions $\varepsilon_{\rm Pointing}$, $\varepsilon_{\rm S_{2}}$, and $\varepsilon_{\rm S_{3}}$ shown inset. When the detector spacing is larger than $20^{\circ}$, the SNN polarization prediction fails. Below a detector spacing of $20^{\circ}$, the model separates input ${\rm S}_{2}$ and ${\rm S}_{3}$ polarizations. When the detector spacing is $\sim 4^{\circ}$, we observe the highest values for ${C_{\rm R}}$. These transitions relate to the fringes produced by PANI-HS shown in Fig. 6(c-f).
In other words, the transitions in the SNN prediction errors are marked by the physical behavior of the film; the compression ratio is highest when the detectors are spaced to sample the interferometric fringe and ${\rm S}_{2}$ polarization modulation at twice the angular frequency or the Nyquist sampling criterion. The encoder has simply mapped pointing and polarization into a angular-varying intensities. Regimes of undersampling and oversampling are identified above and below these transitions. The trends are similar across different SNN models and data preprocessing choices.
@@ Discussion
The ability to encode full-Stokes polarization and achieve polarimetric compressive sensing is a novel implementation of physical encoding with back-end computing. In the past, the encoder algorithmic function is programmed with digital arrays of micromirrors 28, 29, pixelated spatial-light modulator generated masks 30, 31, and printed polarizer arrays or masks 32. While these technologies are well established, most digitally-controlled technologies operate on the principles of filtering, reflection, or refraction, which alone does not spatially separate RCP, LCP, and ±$\pm 45$LP components (i.e., ${\rm S}_{2}$ and ${\rm S}_{3}$).
To spatially separate beams with different polarization ellipticities, the interaction between polarization and pointing or an optical spin-orbit interaction is needed 33, 34. Optical spin-orbit interactions are conventionally operationalized in metasurfaces with a top-down fabrication paradigms and are therefore composed of periodic, high-refractive-index or metal nanostructures 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45. Their order and periodic structures, while theoretically tractable for encoder design, do not provide the wavefront distortion and signal mixing necessary for the compressive sensing of pointing and polarization as illustrated in Fig. 6(a-b).
While random media provides the incoherence function for compressive sensing, too much disorder is difficult to decode. For example, the wave mixing of multimode fiber and multiple-scattering materials is enhanced in longer fiber or with longer optical path lengths in which the decoding algorithms are more involved or less accurate 46, 47, 24, 48, 49, 50, 51, 52. In our work, optical neural networks with random encoders achieve more compression than systems with ordered structures, however random encoders also result in a higher variance in their sensing performance [Fig. 7(c-d)]. Even so, we show that the neural network learning transitions are predicted by the physical characteristics of the metaphotonic thin film. This means that we can, with physical understanding of a thin, self-assembled metaphotonic structure, design for optimal sampling and performance.
PANI-HS is a solution-processed metasurface 23 that spatially separates polarization components and whose thin-layer meso-order is easily decoded with a simple SNN 53. The self-assembled hollow-sphere diffractive films are also lightweight and low-cost, and their fabrication does not require expensive top-down techniques. The primary limitation of PANI is that it may degrade above temperatures of 140 C 54, however other shell materials may be explored in lieu of PANI.
@@ 3 Conclusion
In conclusion, we have demonstrated facile polarimetric compressed sensing with random self-assembled hollow-sphere encoders. An intermediate regime between order (lattices) and disorder (incoherence) is necessary to encode for polarimetric compressed sensing. In PANI-HS, the polarimetric compressed sensing is achieved by three primary responses: an effective birefringence via phase retardance that modulates the diffracted polarization via lattice effects; an effective thin-film interference that modulates the transmitted intensity patterns for pointing via refractive index contrast; and the broadening of the otherwise-sharp angular $k-$space diffracted features that provide incoherence or wavefront distortion via disorder or blur. Our physical encoder functions as a metasurface and a random scattering film, and offers fast and simple SNN decoding for rapid operations. Moreover, we have demonstrated that knowledge of the metaphotonic encoding ( interference and polarization modulation) translates to best sampling for optimal performance.
Hollow encoders produced by templated colloidal assembly represent some of the simplest air-filled lattice designs that serve as a solution-processed metasurface. Our research on templated, self-assembled, hollow encoders not only offer biospeculation about the marvellous vision systems of insects, but also frames a new class of materials for optical sensing and imaging. Our results with a solution-processed and mostly-air metaphotonic film presents new avenues for inexpensive, encoded large-area aperture and windows, for high-dimensional encoded computational cameras, including lightweight, high-speed, low-power, drone cameras and other polarization-sensitive edge computing applications.
@@ 4 Methods
Authors gratefully acknowledge: funding from DARPA YFA 19AP00036 and NSF DMR 1921034 and assistance fabricating samples from S. Weiss.
@@ 5.1 Pair correlation
Figure 8: Optical images used to calculate the order parameters.
@@
To further quantify the degree of ordering by using this synthesis approach, we analyze the distribution of PANI-HS in three randomly chosen areas, and calculate their 2D pair correlation function, $g(r)$ 15, 16,
@@
where a is the shell area, $dn$ is the number of holes that lie within a shell, $r$ is the distance from an arbitrary origin, and $dr$ is the shell thickness. The statistical average of holes is normalized by the average particle number density $\rho\right>$ and the sampling area $da=2\pi rdr$. We analyzed optical images of 900*900 pixels (Figure S1.1) with $r$ = 11 and $dr$ = 0.725. The pair correlation function of PANI-HS and a perfect lattice is plotted in Figure 2a in the main text.
Figure 2a in the main text also shows the first peak of the Fourier transform of each $g(r)$. We compare the full width at half maximum (FWHM), $\kappa$, for the first peak of $g(r)-1$ to that of a perfect lattice ($\kappa_{0}$). The ratio of $\kappa/\kappa_{0}$ is used to quantitatively determine the ordering of the photonic structure, and average from the three areas. A structure with $\kappa/\kappa_{0}\leq 1.5$ is considered very highly ordered. The averaged $\kappa/\kappa_{0}$ for the ordered and random PANI-HS are 1.10 and 2.49 with standard deviation of 0.07 and 0.26.
@@ 5.2 Mesoscale Grain Size Determination
To determine Linear Intercerpt Density (LID) 14, 10000 chords are constructed via the choice of two random, not identical, points in the grain map. These chords are rasterized into a number of pixels, $p_{i}$, along the $i^{t}h$ chord. One of the ends of the chord is chosen as the start; from there, the grain orientation at that pixel is noted and the grain orientation of the next pixel is read. If the grain orientation of the next pixel is sufficiently different than the orientation of the previous pixel (>±.$\pm.5^{o}> ± .$), then $1$ is added to the boundary counter, $b_{i}$. The average grain density is determined by
@@ 6.1 Diffractometry Measurements
Mueller matrix diffractometry measurements are performed using Mueller matrix microscope equipped with a Bertrand lens. The sample is illuminated with a collimated beam, which is collected by a high NA microscope objective. The Bertrand lens focuses the back focal plane of the objective onto the camera, so that it is possible to image the angle-resolved diffraction of the colloidal samples. This instrument uses a computer controlled polarization state generator and a polarization state analyzer to probe the sample with different polarization states, allowing the automatized measurement of the 16 Mueller matrix elements. Further details of the instrument used are available in a prior publication 20.
@@ 6.2 Linear-to-Circular Polarization Conversion
The full transformation from linear-to-circular polarization is observed with non-normal angles of incidence when light is composed of both TE and TM polarizations at an 8.5-degree angle of incidence. With the ordering ratio of $\kappa/\kappa_{0}\leq 1.5$, we observe circular-to-linear polarization conversion as well as linear-to-circular, and note that other polarizations are tuned with an angle of incidence in a manner analogous to metasurfaces that enable full-Stokes transformation and polarization tuning. [See Fig. 11]
@@ 7.1 The dataset
Figure 13: Stokes measurements of disordered sample at higher power for input light of horizontal, vertical, ±$\pm 45^{\circ}$, $RCP$, and $LCP$ polarizations. Both input angle and output angle (with respect to sample normal) vary between $-75^{\circ}$ and $75^{\circ}$.

### v2

@@ 2 Results
When samples are ordered, the simulated resonances are narrow and located at an angle of incidence $\theta_{i}\approx\pm 15-20^{\circ}$. Strong TE modes are observed in ordered PANI-HS samples and disappear as samples are more random, evident from the specular transmission and reflection measurements with a lock-amplifier (See Supporting Information Measurements and Simulations, Fig. S7). In contrast, weak TM-polarized modes in the specular transmission and reflection vary in strength and occur at different incident angles across different experimental batch samples. These variations could relate to the sensitivity of the TM modes to small variations in the sample, i.e., grain mesostructure, fiber nucleation statistics, or PANI doping. However, over many batches and over years in ambient conditions, the sample TE lattice modes $\theta_{i}\approx\pm 15-20^{\circ}$ remain strong and consistent from batch to batch. The periodic modulations of ${\rm S}_{2}$ associated with the TE-polarized lattice modes occur with angular periods of $\sim 40^{\circ}$ and encode the incident state of polarization.
@@
Figure 7: Multi-parameter estimation of incident beam polarization and pointing. (a) Illustration of 4-encoder-detector system for simulation with 150-degree field-of-view. The detector separation for this case is $\Delta\theta_{o}=37.5^{\circ}$. (b) Colormaps illustrating difference LP-signals given the geometry in (a) for raw experimental data of ordered sample (left) and 5-degree window-blurred sample (right). Compression ratios as a function of detector spacing with (c) raw data and (d) blurred data for ordered (blue) and random (green) samples. The markers show the mean values and shaded areas mark the upper and lower bounds from 15 different test simulations with polynomial fits between points. Inset: the pointing and polarization (${\rm S}_{2},{\rm S}_{3}$) prediction errors. Dotted vertical lines mark Nyquist sampling of the pointing or when $\Delta\theta_{o}=$ Pointing MAE $\approx 20^{\circ}$. Best compression ratios are achieved with Nyquist sampling twice the frequency of the interference fringes or $\Delta\theta_{o}\approx 4^{\circ}$.
@@
From this dataset that includes the full-Stokes characteristics at each angle, we consider the sparse “detector measurements” from two crossed LP detectors aligned ±$\pm 45^{\circ}$ from the sensing-plane. Figure 7(b) shows colormaps of the LP-difference detected signals from each encoder given the geometry in Fig. 7(a) for RCP and LCP. The LP-difference detected signals are the logarithms of the intensity signals multipled by the ${\rm S}_{2}$ measurements. The raw experimental data from ordered samples appear noisy and have narrow, angularly-sharp features; in fact, a significant amount of the information lies in the low-power measurements that are noisy. We smooth measurements with a Savitzky-Golay filter [Fig. 7(b,right)]. This filter is analogous to the signal integration that is expected to be present in nocturnal insects and under low-light conditions 25, 26 or with micro-saccadic head movements.
We define the compression ratio ${C_{\rm R}}$ as the improvement of the prediction error with encoding compared to the prediction error from conventional sampling 6. The SNN predictions indicate two notable transitions, which are explained by the encoder characteristics. Figure 7(c-d) shows the compression ratio ${C_{\rm R}}$ where the range of the error predictions $\varepsilon_{\rm Pointing}$, $\varepsilon_{\rm S_{2}}$, and $\varepsilon_{\rm S_{3}}$ are inset. When the detector spacing is larger than $20^{\circ}$, the SNN polarization prediction fails. Below a detector spacing of $20^{\circ}$, the model separates input ${\rm S}_{2}$ and ${\rm S}_{3}$ polarizations. When the detector spacing is $\sim 4^{\circ}$, we observe the highest values for ${C_{\rm R}}$. These transitions relate to the fringes produced by PANI-HS shown in Fig. 6(c-f).
In other words, the transitions in the SNN prediction errors are marked by the physical behavior of the film; the compression ratio is highest when the detectors sample the interferometric fringe and ${\rm S}_{2}$ polarization modulation at twice the angular frequency or at the Nyquist sampling criterion. The role of the encoder is to map the pointing and polarization ellipticity into angular-varying intensity patterns that are, in this case, captured by crossed linear polarizers. Regimes of undersampling and oversampling are identified above and below these transitions. The trends are similar across different SNN models and data preprocessing choices.
@@
The ideal degree of compression occurs for a relatively sparse separation of sensors (2-5) over a wide field of view (±$\pm 75^{\circ}$). Since the ideal sampling depends on the thin-film interference fringes, thicker encoders (larger-diameter hollow spheres or multiple-layer geometries) may be used to decrease the separation of sensors. It is expected, however, that such a change in the encoder will result in trade-offs, i.e., either a narrower field of view or more complicated back-end NNs. Nevertheless, our results suggest that with the compression ratios demonstrated here one could achieve megapixel camera resolution with a kilopixel sensor array, particularly with further optimization with the NN architecture and the PANI-HS nanostructure.
@@ Discussion
The ability to encode full-Stokes polarization and achieve polarimetric compressive sensing is implemented with the back-end computing and the physical encoding from a self-assembled film. In the past, the encoder algorithmic function is programmed with digital arrays of micromirrors 27, 28, pixelated spatial-light modulators 29, 30, and printed polarizer arrays or masks 31. While these technologies are well established, most digitally-controlled technologies operate on the principles of filtering, reflection, or refraction, which alone cannot spatially separate RCP, LCP, and ±$\pm 45$LP components (i.e., ${\rm S}_{2}$ and ${\rm S}_{3}$).
To spatially separate beams with different polarization ellipticities, the interaction between polarization and pointing or an optical spin-orbit interaction is needed 32, 33. Optical spin-orbit interactions are conventionally operationalized in metasurfaces with top-down fabrication paradigms and are therefore composed of periodic, high-refractive-index or metal nanostructures 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44. Their ordered and periodic structures, while theoretically tractable ı.e., straightforward to simulate and analyze with periodic boundary conditions, do not provide the wavefront distortion and signal mixing necessary for compressed sensing. Notably, it is the “imperfections” associated with self assembly (i.e., randomly-oriented nanofibers and sample grain disorder) that provide the structure for physical encoding by which compressive sensing of pointing and polarization is achieved. Moreover and to a large extent, we observe that random structures achieve higher degrees of compression than ordered structures [Fig. 6(a-b)].
However, while random media provides the incoherence function for compressive sensing, too much disorder is difficult to decode. For example, the wave mixing of multimode fiber and multiple-scattering materials is enhanced in longer fiber or with longer optical path lengths in which the decoding algorithms are more involved or less accurate 45, 46, 23, 47, 48, 49, 50, 51. In our work, random encoders also result in less reliable predictions, i.e., a higher variance in their sensing performance [Figs. 7(c-d)]. Even so, we show that the neural network learning transitions are predicted by the physical characteristics of the metaphotonic thin film. This means that we can, with physical understanding of a thin, self-assembled metaphotonic structure, design for optimal sampling and performance.
PANI-HS is a solution-processed metasurface 22 that spatially separates polarization components and whose thin-layer meso-order is easily decoded with a simple SNN 52. The self-assembled hollow-sphere diffractive films are also lightweight and low-cost, and their fabrication does not require expensive top-down techniques. The primary limitation of PANI is that it may degrade above temperatures of 140 C 54, however other shell materials may be explored in lieu of PANI.
@@ 3 Conclusion
In conclusion, we have demonstrated facile polarimetric compressed sensing with random self-assembled hollow-sphere encoders. An intermediate regime between order (lattices) and disorder (incoherence) is necessary to encode for polarimetric compressed sensing. In PANI-HS, the polarimetric compressed sensing is achieved by three primary responses: an effective birefringence via retardance that modulates the diffracted polarization via lattice effects; an effective thin-film interference that modulates the transmitted intensity patterns for pointing via refractive index contrast; and the broadening of the otherwise-sharp angular $k-$space diffracted features that provide incoherence or wavefront distortion via disorder or blur. Our physical encoder functions as a metasurface and a random scattering film, and offers fast and simple SNN decoding for rapid operations. Moreover, we have demonstrated that knowledge of the metaphotonic encoding (interference and polarization modulation) translates to best sampling for optimal performance.
Hollow encoders produced by templated colloidal assembly represent some of the simplest air-filled lattice designs that serve as a solution-processed metasurface. Our research on templated, hollow encoders not only offers insight about the vision systems of insects, but also frames a class of self-assembled materials for optical sensing and imaging. Our results with a solution-processed and mostly-air metaphotonic film open avenues for inexpensive, encoded large-area aperture and windows, for high-dimensional encoded computational cameras, including lightweight, high-speed, low-power, drone cameras and other polarization-sensitive edge computing applications.
@@ 4 Methods
Authors gratefully acknowledge funding from DARPA YFA 19AP00036 and NSF DMR 1921034, which supported all experiments and analyses with exception of the COMSOL simulations and Mueller matrix diffractometry. OA acknowledges TED2021-129639B-I00, CNS2022-136051 (Ministerio de Ciencia, Innovación y Universidades). S. Weiss fabricated the samples sent to OA.
@@ 5.1 Pair correlation
Figure 8: Optical images used to calculate the order parameters. All scale bars represent 5 ${\mu}m$.
@@
To further quantify the degree of ordering by using this synthesis approach, we analyze the distribution of PANI-HS in three randomly chosen areas, and calculate their 2D pair correlation function, $g(r)$ 14, 15,
@@
where a is the shell area, $dn$ is the number of hollow nanospheres that lie within a shell, $r$ is the distance from an arbitrary origin, and $dr$ is the shell thickness. The statistical average of hollow nanospheres is normalized by the average particle number density $\rho\right>$ and the sampling area $da=2\pi rdr$. We analyzed optical images of 900*900 pixels (Figure S1.1) with $r$ = 11 and $dr$ = 0.725. The pair correlation function of PANI-HS and a perfect lattice is plotted in Figure 2a in the main text.
Figure 2e in the main text also shows the first peak of the Fourier transform of each $g(r)$. We compare the full width at half maximum (FWHM), $\kappa$, for the first peak of $g(r)-1$ to that of a perfect lattice ($\kappa_{0}$). The ratio of $\kappa/\kappa_{0}$ is used to quantitatively determine the ordering of the photonic structure, and average from the three areas. A structure with $\kappa/\kappa_{0}\leq 1.5$ is considered very highly ordered. The averaged $\kappa/\kappa_{0}$ for the ordered and random PANI-HS are 1.10 and 2.49 with standard deviation of 0.07 and 0.26.
@@ 5.2 Mesoscale Grain Size Determination
To determine Linear Intercerpt Density (LID) 53, 10000 chords are constructed via the choice of two random, not identical, points in the grain map. These chords are rasterized into a number of pixels, $p_{i}$, along the $i^{th}$ chord. One of the ends of the chord is chosen as the start; from there, the grain orientation at that pixel is noted and the grain orientation of the next pixel is read. If the grain orientation of the next pixel is sufficiently different than the orientation of the previous pixel (>±.$\pm.5^{o}> ± .$), then $1$ is added to the boundary counter, $b_{i}$. The average grain density is determined by
@@ 6.1 Diffractometry measurements
Mueller matrix diffractometry measurements are performed using Mueller matrix microscope equipped with a Bertrand lens. The sample is illuminated with a collimated beam, which is collected by a high NA microscope objective. The Bertrand lens focuses the back focal plane of the objective onto the camera, so that it is possible to image the angle-resolved diffraction of the colloidal samples. This instrument uses a computer controlled polarization state generator and a polarization state analyzer to probe the sample with different polarization states, allowing the automatized measurement of the 16 Mueller matrix elements. Further details of the instrument used are available in a prior publication 19.
@@
6.2 Conversion of Jones to Stokes polarization measures
We convert our model from Jones to Stokes vectors since each representation has different advantages. The input and diffracted Jones vectors $\mathbf{E}_{in}=E_{o}[A_{x},A_{y}\exp(-i\delta)]^{T}$ are shown with Stokes measures ${\rm S}_{1}=A_{x}^{2}-A_{y}^{2}$, ${\rm S}_{2}=2A_{x}A_{y}\cos(\delta)$, and ${\rm S}_{3}=2A_{x}A_{y}\sin(\delta)$, where $A_{x}^{2}+A_{y}^{2}=1$. Additionally, $\delta=\pm\pi/2$ and ${\rm S}_{3}\pm 1$ for RCP/LCP, respectively.
@@ 6.3 Linear-to-Circular Polarization Conversion
The full transformation from linear-to-circular polarization is observed with non-normal angles of incidence when light is composed of both TE and TM polarizations at an 8.5-degree angle of incidence. With the ordering ratio of $\kappa/\kappa_{0}\leq 1.5$, we observe circular-to-linear polarization conversion as well as linear-to-circular, and note that other polarizations are tuned with an angle of incidence in a manner analogous to metasurfaces that enable full-Stokes transformation and polarization tuning [See Fig. 11].
@@ 7.1 The dataset
Figure 13: Stokes measurements of disordered sample at higher power for input light of horizontal, vertical, ±$\pm 45^{\circ}$, ${\rm RCP}$, and ${\rm LCP}$ polarizations. Both input angle and output angle (with respect to sample normal) vary between $-75^{\circ}$ and $75^{\circ}$.